OpenStudy (anonymous):

Simplify the radical: square root of 50

OpenStudy (anonymous):

25 * 2 = 50 so take the square root of 25 which is 5 out of the radical and leave the 2 5sqrt2

OpenStudy (anonymous):

\[\sqrt{50}=\sqrt{2\times5\times5}=\sqrt{2\times5^{2}}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure :)